It is not safe to put aerosol canisters in a campfire, because the pressure inside the canisters gets very high and they explode. If Cindy has a 1.0 L canister that holds 2 moles of gas, and the campfire temperature is 1400°C, what is the pressure of the canister in kPa? 


1
Expert's answer
2021-05-03T07:58:29-0400

PV = nRT

 

P = pressure

 

V = volume = 1.00 L = 0.00100 m3

 

n = 2.00 moles

 

T = absolute temperate = (1400 + 273) K = 1673 K

 

R = universal gas constant = 8.31 J / (mole K)

 

P = nRT / V = (2.00 moles) [8.31 J/(mole K)] (1673 K) / (0.00100 m3)

 

P ≅ 2.78 x 107 N/m2